Aos? sample of supply, what the empirical formula Ho is of hydrocarbon? this C H and o) is burnes What is the name formula the same as the empirical formula of Write down the equation of the combustion reaction for this hydrocarbon and balance it. SHOW YOUR WORKI (keep mal in points 11) If 40 g of water are heated from 22.cto 46 c, how much heat in joules is gained by the water given that cs for water is 4.184 J/e c sHow YoUR wORKI (10 Pts) 12)A photon has an energy (E) 2.76 x 10 19 J. What is its wavelength? (A Pts) [Plancks constant is 6.626x 10 34 Js] what color is this photon? (1 Pt
media%2Fdf4%2Fdf4dac92-0d26-4b1e-959d-07
media%2F644%2F64460324-3f89-4863-be3f-7d
media%2Fe00%2Fe00732b1-00a6-4cea-a8bf-8a
Answer #1

Question 10.

m = 0.57 g of CHO

m = 1.1 g of CO2, mol = mass/MW = 1.1/44 = 0.025 mol of C

m = 0.65 g of H2O, mol = masS/MW = 0.65/18 = 0.03611 mol of H2O --> 0.07222 mol of H

then

mass of C = 0.025 *12 = 0.3 g

mass of H = 0.07222 *1 = 0.07222 g

mass of O = 0.57 - 0.3 - 0.07222 = 0.19778 g of O

mol = mass/MW = 0.19778/16 = 0.01236 mol of O

ratios

0.025 /0.01236 = 2

0.07222 /0.01236 = 5.8

ratio is then

C2H6O ; this is most likely ethanol, CH3CH2-OH

The equatio of combustion

C2H6O + O2 = CO2 + H2O

balance

C2H6O + 3O2 = 2CO2 + 3H2O